### v3

# Local Identification in Instrumental Variable Multivariate Quantile Regression Models

Abstract
In the instrumental variable quantile regression (IVQR) model of [chernozhukov2005iv], a one-dimensional unobserved rank variable monotonically determines a single potential outcome. Even when multiple outcomes are simultaneously of interest, it is common to apply the IVQR model to each of them separately. This practice implicitly assumes that the rank variable of each regression model affects only the corresponding outcome and does not affect the other outcomes. In reality, however, it is often the case that all rank variables together determine the outcomes, which leads to a systematic correlation between the outcomes. To deal with this, we propose a nonlinear IV model that allows for multivariate unobserved heterogeneity, each of which is considered as a rank variable for an observed outcome. We show that the structural function of our model is locally identified under the assumption that the IV and the treatment variable are sufficiently positively correlated.
Keywords— IVQR, local identification, multivariate quantiles, optimal transport
1 Introduction
The estimation of quantiles of potential outcomes in the presence of endogeneity has received increasing attention. For this purpose, the instrumental variable quantile regression (IVQR) model introduced by [chernozhukov2005iv] is popular in applications. While the IVQR model assumes that outcome variables are one-dimensional, researchers are often interested in multiple outcomes at the same time. For example, there is much work that studies Engel curves to investigate how the shares of food and leisure expenditures change as the total expenditure changes using the gross earnings of the head of household as an instrument. See for example [blundell2007semi], and [imbens2009identification]. In such cases, researchers typically run IV regression models separately for each dimension of the outcome variable. However, this approach can be problematic for several reasons.
First of all, it ignores how different outcome variables are correlated with each other, and consequently it is silent on the joint distribution of them. In the Engel curve example, the expenditure shares of food and leisure are likely correlated, not only because there is substitutability, i.e., more food, less leisure, but also because their sum must not exceed one.
More importantly, applying the standard quantile model separately makes little sense in many cases because one of the crucial assumptions of [chernozhukov2005iv], the rank similarity, is unrealistic. Specifically, in order to apply [chernozhukov2005iv] to the Engel curve dataset, it must be assumed that individuals choose the expenditure share on food based only on their preference for food. However, this is unlikely to be true, because they should also care about their preference for leisure, which determines how much they can spend on food.
To address these issues, we propose a novel nonlinear model that allows for multidimensional outcome variables that are potentially correlated with each other. We consider a multivariate extension of the potential outcome framework of [chernozhukov2005iv]. For a treatment let be the corresponding -dimensional potential outcome. We assume that conditional on covariates, it is represented as where is a structural function and is a -dimensional rank variable. The rank variable is a normalized random variable and accounts for the heterogeneity of potential outcomes among observationally identical individuals. It can be interpreted as representing unobserved individual characteristics.
Existing work on endogenous quantile models imposes various monotonicity assumptions on the structural function for identification. Generalizing the assumption of [chernozhukov2005iv] that the structural function is the quantile function of the potential outcome, we assume that the derivative of the structural function is symmetric and positive definite. Under this restriction, an increase in the th entry of leads to an increase in the th entry of In addition, an increase in a particular entry of can affect other dimensions of both positively and negatively. This allows the model to capture substitutability/complementarity between different dimensions of the outcome. Running the standard IVQR separately does not have this flexibility.
A vector-valued function with a symmetric and positive definite derivative is characterized by the cyclical monotonicity, which is a notion of monotonicity for vector-valued functions. There are many economically important examples of cyclically monotone functions in contract theory ([rochet1987necessary], [kushnir2021monotone]), revealed preference theory ([ekeland2013housing]), and econometrics ([shi2018estimating], [chernozhukov2021identification]). We show in Section 2.2 that under a reasonable setup, the vector of the optimal expenditure shares of goods is written as such.
In the optimal transport literature, a vector-valued function with a symmetric and positive definite derivative is known to minimize the quadratic transportation cost and called an optimal transport map. Moreover, recent developments in statistical theory have revealed that it has many desirable properties that qualify it as an extension of the quantile function to random vectors (see, e.g., [ekeland2012comonotonic], [chernozhukov2017monge], [hallin2021distribution], [ghosal2022multivariate]). This justifies calling our model an IV multivariate quantile regression model.
Regarding identification, [chernozhukov2005iv] show that in order to identify quantiles of one-dimensional potential outcomes for binary treatments, an IV with binary support satisfying what they call the full rank condition is necessary. For -dimensional potential outcomes, it may seem natural that the support size of an IV needed for the identification is greater than two because there are functions to be identified. Surprisingly, however, our main result, Theorem 3.1, states that an IV with only binary support can achieve the desired identification. At the cost of multidimensionality, the IV is required to be correlated with the treatment status sufficiently positively, which is a stronger condition than the full rank condition, and it becomes more demanding as the dimension of the potential outcomes increases.
Related literature. Our model stands on the IVQR model that is proposed and developed by [chernozhukov2005iv], [chernozhukov2006instrumental], [chernozhukov2007instrumental] and [chernozhukov2013quantile]. Unlike these work, we allow for the multidimensionality in the outcome and unobservable heterogeneity. There are different studies that deal with endogeneity when outcome variables are one-dimensional. [abadie2002instrumental] consider quantile treatment effects based on the local average treatment effect framework developed by [imbens1994identification]. Their approach identifies treatment effects on compliers and is valid only for binary treatments. Extending [chesher2003identification], [imbens2009identification] take the control function approach under the assumption that the selection equation is monotone and the instrumental variable is independent of unobservable disturbances. [gunsilius2023condition] extends the control function approach to the multidimensional outcome setting. [matzkin2008identification] considers nonparametric identification in simultaneous equation models. In the paper, endogenous variables are assumed to be continuous, and the number of them must be equal to the dimension of unobserved heterogeneity, whereas our model allows for discrete treatments and arbitrary randomness in the treatment assignment stage. In the simultaneous equation framework, [blundell2017individual] propose a nonseparable model when a proxy variable of unobservable heterogeneity is available.
Our assumption on the structural function requires it to be a multivariate quantile function that is defined and developed by [ekeland2012comonotonic], [chernozhukov2017monge], [hallin2021distribution], and [ghosal2022multivariate]. Multivariate quantile functions are defined based on the optimal transport theory ([villani2003topics], [villani2009optimal]). [galichon2018optimal] provides a thorough review on economic applications of optimal transportation. Mathematical results from optimal transport theory are used to prove the identification.
Organization: This paper is organized as follows. Section 2 introduces the IV multivariate quantile regression model. Section 3 formally states the local identification result for the model. Section 4 discusses the cost of taking multidimensional potential outcomes into account and outlines the proof of our main theorem. All proofs and mathematical preliminaries are provided in the Appendix.
2 IV multivariate quantile regression model
2.1 Model
Let be a -dimensional potential outcome vector, an endogenous variable (treatment), an exogenous covariate vector, and an IV. For now, the treatment and IV can be either discrete or continuous, although we will focus on the discrete cases for identification in the next section. Let be a compact convex set with a piecewisely boundary Let be a reference probability measure of which support is Assume is absolutely continuous with respect to the -dimensional Lebesgue measure.
Our model extends [chernozhukov2005iv] so that it accounts for multidimensional outcome vectors. The following is the primitives.
Assumption 2.1.
The random variables satisfy the following conditions with probability one:
For each it holds that and there exists a function such that is continuously differentiable in the first argument and its derivative is symmetric and positive definite on
For each is independent of conditional on 111This is slightly redundant. It is enough to assume that the conditional independence between and for some but we keep the stronger assumption following the literature.
for some unknown function and random element
Conditional on are identically distributed.
The observed random variables consist of and
We call this model an IV multivariate quantile regression model. The use of “multivariate quantile” follows the literature on the extension of classical quantile functions to multivariate random variables. Section 2.4 discusses this point in more detail.
The difference from [chernozhukov2005iv] appears in (A1). It follows from Theorem A.1 that for any reference measure there exists with a symmetric and positive semi-definite derivative such that but it is not necessarily regular as specified in (A1). Assumption (A1) assumes the smoothness of and the strict positive definiteness of its derivative. This assumption holds if both and the distribution of conditional on are sufficiently regular, according to the regularity theory of optimal transport (see, for example, [caffarelli1992regularity]). Hence, for any other sufficiently regular reference measure condition (A1) is satisfied with some and This implies that any causal parameter defined as a functional of the distributions of the potential outcomes, such as average/quantile treatment effects, is independent of the choice of the reference measure. Also, the continuity of imposed in (A1) implies that the range of is compact and so is the support of conditional on
Condition (A2) implies that the IV is independent of potential outcomes conditional on covariates. Condition (A3) is a weak restriction that allows a broad class of assignment rules of treatments. In particular, in (A3) can be correlated with potential outcomes. Condition (A4) is called the rank similarity, which is usually assumed in the literature. This is weaker than the rank invariance assumption that does not change with assigned treatments. For more details of these assumptions, see [chernozhukov2005iv] and [chernozhukov2013quantile].
The unobservable random variable accounts for the heterogeneity of outcomes among the observationally identical individuals. It is interpreted as a rank variable in the sense that the th entry of is monotonically determined by the th entry of conditional on covariates Indeed, by the positive definiteness of the (Assumption (A1)), it holds that where and On top of that, unlike the one-dimensional model, it is also possible that the th entry of affects the th entry of for in a flexible way. More precisely, the partial derivative can be either positive or negative as long as is positive definite, which allows for the correlation between different dimensions of the observed outcome vector.
Technically, one can choose any absolutely continuous probability measure on any compact convex as the reference measure but taking and seems natural in the examples considered in this paper. One could also think of the -dimensional unit ball as and the spherical uniform distribution as as considered in [chernozhukov2017monge] and [hallin2021distribution].
By condition (A1), the observed outcome is written as the structural form where The following representation is the main testable implication of the model.
Theorem 2.1.
Suppose Assumption 2.1 holds. Then, it holds with probability one that for each measurable
where In particular, holds.
As the LHS of (1) is determined by the joint distribution of the observable variables, the equation gives a conditional moment restriction. That is, any candidate structural function necessarily satisfies (1) in which is replaced by For identification, we consider conditions under which is the only functions that satisfy the equation in the latter section.
For the notation simplicity, let by suppressing covariates and all the following analysis should be understood as conditional on This treatment does not affect identification results because the distribution of observables conditional on is identifiable.
When outcomes are univariate and the reference distribution is the uniform distribution on the unit interval, as in [chernozhukov2005iv], is the -quantile of and satisfies
for all which corresponds to setting and in (1). If the treatment variable is supported on a finite set , this is a simultaneous equation with unknown variables Hence, if the instrument takes on more than or equal to values and the equation system is non-degenerate, it is expected that the solution is unique, at least locally. Indeed, [chernozhukov2005iv] show that this observation is correct. Checking the condition for each quantile level establishes the nonparametric identification of the structural function
This approach works in dimension one because for each the values are determined in the conditional moment restriction (1) by choosing a specific set Unfortunately, it does not apply to higher dimensions. To see this, for consider and fix Then, any set does not make (1) an equation of only, because the image can be fully nonlinear. Specifically, the image set depends on the values of not only at but also at all the boundary points of Since the set is an infinite-dimensional object, considering (1) for fixed separately does not provide any identification power with a finitely supported IV. To overcome this difficulty, we instead think of (1) as a measure-valued equation in order to exploit the relationship between and for different sets which is not necessary for the identification of univariate structural functions.
2.2 Example: the Engel curve
We illustrate an example the multivariate IV quantile model naturally fits. We consider the Engel curve and investigate how the shares of expenditures on goods change as the total expenditure varies using the gross earnings of the head of household as an instrument.
Suppose we are interested in two goods with price We combine all the other goods into one and call it that has price These prices are assumed to be fixed. A consumer earns He decides his consumption schedule His utility from consuming and paying is where is a concave function and represents his “taste” about the goods. The first two terms are obtained by decomposing a concave utility function into the linear part and the other This formulation assumes that consumers are heterogeneous in the coefficients of the linear component of the utility function. He can choose to save instead of consuming. By saving he receives a (discounted) value Thus, he solves the following maximization problem:
where This is rewritten as
where the new variable represents the total expenditure. This formulation reflects the fact that people typically determine how much to save first and then allocate the budget to the goods. Let be the maximizer of the outer maximization problem and be that of the inner one at For fixed expenditure level we denote the optimal expenditure share of good by
As the following proposition claims, the optimal share depends only on a transformed preference measure where for The vector can be thought of as a relative propensity to consume compared to other goods in the consumption basket.
Proposition 2.1.
Assume and Then, given total expenditure level the optimal expenditure share vector is a function of and written as where is a convex function In particular, if is differentiable, its derivative is symmetric and positive semi-definite.
Suppose that consumers are randomly given the gross earnings and the preference vector which are assumed to be independent. They maximize (2) and find the optimal total expenditure and shares where is associated with If an econometrician observes but not and then this setup is consistent with the IV multivariate quantile regression model.
2.3 Comparison with [chernozhukov2005iv]
Even when several outcome variables are present, the one-dimensional quantile model of [chernozhukov2005iv] could be applied to each dimension. However, this approach has some problems. First of all, it does not account for correlation between different outcome variables. In the Engel curve example in the previous subsection, the expenditure shares of food and leisure are correlated, because their sum must be no more than one.
More importantly, the dimension-wise approach typically leads to the violation of the rank similarity assumption of [chernozhukov2005iv]. For consider a binary treatment environment and assume that satisfies Assumption 2.1 with the rank invariance and where is a random variable independent of all the other variables. The potential outcome for treatment is componentwisely represented as
Since and are independent, it is expected that the treatment effect of is identified if is sufficiently informative.
Consider an empirical researcher who is interested only in the first component of the potential outcome. In the Engel curve example, this describes the situation where the researcher is interested only in the expenditure share of food. In this case, it is common to apply the standard quantile model of [chernozhukov2005iv] to the data To see that this practice leads to the violation of the rank similarity assumption, notice that the structural equation under the one-dimensional model is where and is the quantile function of and that is measurable in via The rank similarity requires but this is satisfied only when which does not hold in general.222For a counterexample, consider and In other words, the instrument is invalid in the sense that it is correlated with the unobserved heterogeneity via the treatment variable
The failure of the rank similarity is caused by the fact that depends not only on but also on where and [chernozhukov2005iv] do not allow such dependence, which is often unrealistic. Specifically, in order to apply [chernozhukov2005iv] to the Engel curve dataset in the previous subsection, it must be assumed that individuals choose the expenditure share on food based only on their preference for food. However, they should also care about their preference for leisure, since it determines how much they can spend on food. Our multivariate quantile model allows for the dependence of the potential outcomes on the multiple unobservables which remedies the problem.
In summary, the IV multivariate quantile model enables to consider multiple outcome variables at the same time, since it captures the correlation between different outcomes. Moreover, even if only some of the outcome variables are of interest, our model can alleviate the endogeneity by considering other outcomes together.
2.4 Connections to optimal transport theory
Our multivariate outcome model is tightly related to the optimal transport theory. By Assumption 2.1, the structural function transports to which is denoted by where is the distribution of that is, it satisfies for all measurable Also, remember that has a symmetric and positive definite derivative. It is well-known that is optimal in the sense that it minimizes the quadratic transportation cost; that is, solves
where is the Euclidean norm, and it is called the optimal transport map from to
It is also known that the optimal transport map characterizes the target distribution once the source distribution is fixed. It also has a certain monotonicity property, called the cyclical monotonicity. Recent developments in statistical theory have shown that the optimal transport map can be thought of as a multivariate extension of quantile functions (see, e.g., [ekeland2012comonotonic], [chernozhukov2017monge], [hallin2021distribution], [ghosal2022multivariate]). In this literature, the optimal transport map from to is also called the multivariate quantile function of with respect to the reference measure Following this terminology, we call our model an IV multivariate quantile regression model.
See Appendix A for more details.
3 Identification
In this section, we consider the identification problem of the structural functions. We assume that the treatment and IV are supported on the same finite set, i.e., We will further restrict our attention to binary treatments/IVs in the later part of this section. More general cases are discussed in Appendix D.
The representation (1) of Theorem 2.1 and the change-of-variables formula imply that the true structural functions solve the functional equations
where
For a scalar function a vector-valued function and a matrix-valued function we define three types of supremum norms as follows:
For a compact set in a Euclidean space, let be the set of -valued functions on that are -times continuously differentiable on and the derivatives can be continuously extended to
Our identification result below shows that is the locally unique solution to (3) in a certain function class. To state the result formally, consider the normed space equipped with Also, for fixed constants define a subset of as
where is the largest eigenvalue of and define similarly. For a constant the space in which the parameter is identified is
Since we are interested in identification, we assume the correct specification.
Assumption 3.1.
Assumption 3.1 is weak under Assumption 2.1, as it just requires that be smooth and that it have moderate derivatives.
The conditions in the definition in restricts possible deviations from the true structural function. This type of condition is called a source condition and usually assumed in the literature of inverse problems and nonparametric IV identification. See, for example, [carrasco2007linear], [chen2014local], [scaillet2016ill], and [centorrino2024iterative]. Section 2.3 of [chen2014local] discusses an example in which identification fails if an analogous condition is dropped.
We further impose two regularity conditions. Let be the support of the distribution of
Assumption 3.2.
For each the support is a convex compact set with a piecewisely boundary
Assumption 3.3.
For each
Assumption 3.2 requires that the support of a potential outcome be regular. The convexity of the support is necessary for many purposes in the optimal transport theory (see, for example, [villani2003topics] and [figalli2017monge]). The smoothness of the boundary is rarely a problem in practice, as it allows for kinks at some points. Combined with Assumption 3.2, Assumption 3.3 implies that is bounded.
We consider the identification of structural functions in the following sense.
Definition 3.1.
Suppose We say the structural function is identified in if the following holds: If a set of random variable satisfies Assumption 2.1 with some then the joint distribution of is different from that of Also, we say the structural function is locally identified in if there exists such that is identified in
In what follows in this section, we focus on the case of i.e., to develop the idea clearly. Identification results for nonbinary treatments are given in Appendix D.
The following is our main theorem.
Theorem 3.1.
Condition (4) is a novel assumption that we cannot find in the literature. Roughly speaking, this condition requires () to be sufficiently positively correlated with (). Too see this, let be the density of conditional on and and suppose that it satisfies for some constants Then, the condition
is sufficient for condition (4). This holds if and are large enough. Since condition (4) requires a positive correlation between and the labeling of the IV should be switched in favor of the condition.
4 Discussions
4.1 Cost of multidimensionality
When condition (4) can be relaxed as follows.
Proposition 4.1.
Condition (4) is stronger than matrix (5) being positive definite, which can be shown as follows. Since equation (4) implies
and therefore, the determinant of matrix (5) is strictly positive. As its trace is clearly positive, it is positive definite.
Proposition 4.1 is consistent with [chernozhukov2005iv], who show that the identification holds if matrix (5) is full-rank, because the determinant of any full-rank matrix can be positive by switching columns if necessary.
The identification for holds under a milder assumption because the proof of Proposition 4.1 uses a fact that is valid only in dimension one. Specifically, it depends on the fact that the cofactor matrix333For an invertible matrix its cofactor matrix is defined as of any matrix is the identity if only if See Lemma C.4 for details. This implies that the positive definiteness of matrix (5) being replaced by condition (4) can be seen as the cost of taking multidimensional potential outcomes into account.
As the dimension increases, condition (4) becomes more demanding. This is quite natural, as one needs to identify higher dimensional structural functions. However, the minimum size of the support of the IV, needed for the identification is independent of the dimension of the potential outcomes, as long as the IV is informative enough.
4.2 Outline of the proof of Theorem 3.1
Roughly speaking, the proof of Theorem 3.1 consists of two steps. In the first step, we show that the support of the potential outcome is identified under condition (4). If a candidate of the structural function induces a different support of the potential outcome, then it cannot be consistent with the joint distribution of the observable variables Thus, we can focus on the identification in the set of structural functions that preserve the support.
In the second step, we show that is the locally unique solution to the system (3) of equations in the set of structural functions. We use an implicit-function-theorem-type argument that is common in the literature on the identification of finite-dimensional parameters in nonlinear models. We consider the linearization of the nonlinear system (3) around the truth It turns out that the positive correlation condition (4) implies that the “slope” of the linearized system is away from zero, which shows the local uniqueness of the solution.
Appendix A Preliminaries
In this section, we briefly review mathematical concepts that are used in the paper. Fix a probability measure on a compact convex subset of the -dimensional Euclidean space, and let be a -dimensional random vector. First, consider the case of and According to the inverse probability integral transform, there exists a random variable such that almost surely where is the quantile function of (see, for example, Proposition 3.2 of [shorack2000probability]). Moreover, is the (almost surely) unique nondecreasing function that satisfies this equation for some uniform random variable. This observation for the classical one-dimensional case holds more generally by replacing the monotonicity of with the restriction that it is written as the gradient of a convex function.
Theorem A.1.
Let be an absolutely continuous probability measure on and be an absolutely continuous -dimensional random vector. There exists a convex function and a random vector such that almost surely and Furthermore, is -almost surely unique.
This is a version of the fundamental theorem in the optimal transport theory known as the Brenier-McCann theorem ([brenier1991polar] and [mccann1995existence]). The gradient is called the optimal transport map from to the distribution of under the quadratic cost because it solves the following Monge problem:
where is the Euclidean norm and is the distribution of
Given the reference probability measure the distribution of is characterized by the gradient via The function is called the multivariate quantile function of The multivariate quantile function has a symmetric and positive semi-definite derivative and therefore, it is cyclically monotone, that is, for any with it holds that
as is shown in [rockafellar1966characterization]. Conversely, if a function has a symmetric and positive definite derivative, then it is written as the gradient of a convex function.
Lemma A.1.
Suppose that a differentiable function has a symmetric and positive definite derivative on Then, there exists a strictly convex differentiable function such that on Furthermore, is unique up to an additive constant.
Furthermore, a multivariate quantile function induces a bijection between the interior of the domain and that of the range under some regularity conditions. For similar results, see Theorem 1 of [cordero2019regularity] and Proposition 3.1 of [ghosal2022multivariate].
Theorem A.2.
Suppose that a continuous function has a symmetric and positive definite derivative on Also, assume that its image is convex. Then, is a continuous bijection between and
Appendix B Local identification for system of equations
B.1 Setup and theorem
The proof of Theorem 3.1 relies on a more general identification result (Theorem B.1), which is of independent interest. We devote this section to the identification result that can be applied to parameters that are characterized by the solution of a system of equations.
Let be normed spaces, be any subset, and be a finite set. For fix and We consider the following system of equations
Suppose is a solution to (7). We are interested in whether is the locally unique solution. To formulate deviations from we define the tangent set at as
The set specifies possible deviations in from Notice that if then We also define the set of normalized tangent vectors as
Remember that when and are finite-dimensional, the locally unique solvability of a system of equations is given by the implicit function theorem, which assumes the differentiability and the full rankness of the Jacobian matrix. As our identification result is an analogue statement for infinite-dimensional parameters, we impose assumptions corresponding to these two.
The first assumption is the differentiability of at
Assumption B.1.
For there exists such that
where if the set over which the supremum is taken is empty, the value reads zero.
It is clear that if Assumption B.1 is reduced to the standard Fréchet differentiability. For strict subsets of it is weaker than the Fréchet differentiability. In particular, even if is not a vector space, Assumption B.1 is well-defined. Unlike the Fréchet differentiability, the linearity of is not required.
The next assumption is an infinite-dimensional version of the full rankness of the Jacobian matrix.
Assumption B.2.
It holds that
Now, we are ready to state the main theorem in this section.
Theorem B.1.
This result is closely related to Section 2 of [chen2014local]. Indeed, their Theorems 1 can be shown from Theorem B.1. Our theorem has a different focus than their result in that it covers the local identification problem for systems of equations and that the domain of functionals is not assumed to be a Banach space.
B.2 Proof of Theorem B.1
We first show an equivalent representation of Assumption B.2.
Lemma B.1.
Let Assumption B.2 holds if and only if there exists such that for all
Proof of Theorem B.1.
Take in Lemma B.1. By Assumption B.1, there is such that for all and
Fix and suppose that with satisfies for We will show By Assumption B.1,
where the equality holds because the first inequality holds because and the last one holds because By Assumption B.2 and Lemma B.1, Therefore,
which immediately implies ∎
Appendix C Proof of Theorem 3.1
Proof of Theorem 3.1.
We first show that the range of any candidate of the structural function must be the same as that of the truth which is
Lemma C.1.
Let Under condition (4), it holds
Since the set in the RHS of the statement is identified, if a candidate has a different range than the joint distribution of in the notation in Definition 3.1, cannot be the same as that of Thus, it is enough to show the local identification in
where
Next, we show that is the locally unique solution to the system (3) of equations in using Theorem B.1. Recall that is a normed space and that Let be the space of finite signed Borel measures on equipped with the total variation norm:
Then, is a normed space. For define as
Then, the identification restriction (3) is written as for all We apply Theorem B.1 to this system of equations where and The normalized tangent set that appears in Appendix B is
We need to check Assumptions B.1 and B.2. Remember that for a smooth vector field the divergence of at is defined as
The following lemma states that is differentiable in the sense of Assumption B.1.
Next, we show the Assumption B.2 by contradiction. Suppose there is a sequence such that
as By definition, there exists such that As and are optimal transport maps, they are written as and for some convex functions and By adding a constant, we may assume without loss of generality. Then satisfies and Since there exists such that it holds that for any
where Hence, we have
where the last inequality holds because
Let As we have which is independent of Also, it holds that and that
By considering a function in the definition (8) of the total variation norm, the integration by parts implies
where
where is the outward normal unit vector of at and is the -dimensional surface measure on
The folloiwng lemma shows that is nonnegative. Remember that since is continuous, implies
Lemma C.3.
For and define Then, at any point is an outward normal vector of
By Lemma C.3, we have
Since the support is convex by Assumption 3.2, it holds that for which implies Thus, we have and therefore,
Taking the sum over yields
Now, we restrict our attention to the case of The following lemma shows the positive definiteness of the integrand of the RHS of (10).
Lemma C.4.
By Lemma C.4, the integrand in the RHS of (10) is almost surely nonngetive. Since the LHS of (10) converges to zero by (9) and we have
for almost all by taking a subsequence if necessary. Recall that and since Then, the sequence is uniformly bounded and uniformly equicontinuous, and therefore, there exists a uniformly converging subsequence by the Arzelà–Ascoli theorem. Let be the limit function. By equation (11), almost everywhere by Lemma C.4, and it holds everywhere, as is continuous. Hence, the uniform convergence implies up to a subsequence. However, this contradicts to the fact that which particularly implies Thus, Assumption B.2 is satisfied.
Appendix D Generalization of Theorem 3.1
In this section, we state a generalization of Theorem 3.1 that allows for nonbinary treatments and IVs. The proof is a straightforward extension of that of Theorem 3.1 once we replace condition (4) with the following.
Assumption D.1.
There is a constant matrix such that for
holds for -almost all
Proof.
The proof is similar to that of Theorem 3.1. Lemma C.1 holds similarly under Assumption D.1, and Assumption B.1 holds as in Lemma C.2. We will verify Assumption B.2. For let
Clearly, Also, let Then, we have and Since
where it holds Notice also that Hence, by considering a function we have
where the last inequality follows by Lemma C.3. Moreover, by and Assumption D.1, it holds
The convergence of the LHS and Assumption D.1 imply that the integrand in the RHS converges to zero almost everywhere on up to a subsequence. Assumption D.1 also implies almost everywhere up to a subsequence. The Arzelà-Ascoli argument concludes which is a contradiction. ∎
Appendix E Omitted proofs
E.1 Proof of Theorem 2.1
E.2 Proof of Proposition 2.1
Proof.
The optimization problem (2) can be written as
where the new variable can be understood as the total expenditure. The inner maximization problem is written as
where and Note that is the budget share of good By eliminating this problem is equivalent to
where and which is concave in Here, we can think of as the consumer’s relative propensity to consume good Under the assumption that this problem has an interior solution, if the consumer has consumption propensity and the total expenditure is the optimal shares of goods are given by
where is the convex conjugate of ([rockafellar1970convex]). Since is convex, the argument above implies that the vector of the optimal expenditure shares has a symmetric and positive semidefinite derivative, if it exists. ∎
E.3 Proof of Proposition 4.1
E.4 Proof of Theorem A.1
Proof.
By the Brenier-McCann theorem, there exists a convex function on such that where is the distribution of Let Then, it holds that Let be the Legendre transform of that is, Clearly, is convex. As for -almost all we have almost surely. The uniqueness of is a direct consequence of the Brenier-McCann theorem. ∎
E.5 Proof of Lemma A.1
Proof.
As is symmetric on which is a convex domain, there exists such that by Poincaré’s lemma. Since is positive definite, is strictly convex. The uniqueness follows immediately from Lemma 2.1 of [del2019central]. ∎
E.6 Proof of Theorem A.2
Recall from the statement of Theorem A.2 that is continuous, surjective, and has a symmetric and positive definite derivative on We first provide three auxiliary lemmas.
Lemma E.1.
The function is strictly cyclically monotone on i.e., it is cyclically monotone, and the inequality of (6) is strict unless
Proof.
Lemma E.2.
The function is injective on
Proof.
Let be a convex function that satisfies the condition in Lemma A.1 for The strict convexity of implies that is injective on because its subdifferentials are disjoint. ∎
Lemma E.3.
For it holds that
Proof.
The proof is by contradiction. Suppose there is such that Since is convex, there exists such that for by the separating hyperplane theorem. It follows that for sufficiently small By Lemma E.1, we have Take a sequence such that Then, a contradiction. ∎
E.7 Proof of Lemma B.1
Proof.
The “if” part holds because for any
To show the “only if” part, suppose
where by the hypothesis. Then, we have
E.8 Proof of Lemma C.1
Proof.
Let such that for some Let Since we have where is the -ball around Thus, we have where the first equality holds by Assumption (A2). This implies and “” holds by taking the closure.
Let and suppose is not in the RHS. Then, there is such that for all and However, this implies that condition (4) fails on the ball, which has a positive measure. Hence, lies in the RHS, and “” holds by taking the closure. ∎
E.9 Proof of Lemma C.2
We first show an auxiliary lemma on the uniform differentiablity.
Lemma E.4.
Let be a subset of an inner product space Suppose the function is (Fréchet) differentiable on the interior of with a uniformly continuous derivative Then, is uniformly differentiable, i.e.,
Proof.
Let Since is uniformly continuous, there is such that
Fix and choose such that and Since
we have
Remark E.1.
In general, the integral of should be understood in the sense of Bochner, but this complication is not important for the proof of Lemma C.2.
For matrices let
For an invertible matrix define its cofactor matrix as
We rewrite the formula in Lemma C.2.
Lemma E.5.
For and it holds
Proof.
For simplicity of notation, we omit arguments of functions.
By the definition of the cofactor matrix, we have
For the second term, it holds
where the last equality follows from Piola’s identity. See, for example, the lemma in page 440 of [evans2010partial]. ∎
Proof of Lemma C.2.
By Lemma E.4, we have
for small because by the definition of and because is continuously differentiable on by Assumption 3.3. Thus, it holds that
for small
Similarly, by Lemma E.4 and Jacobi’s formula, we have
for small because the eigenvalues of lie between and and because is of on This gives a bound for as follows:
for small as by the definition of
Finally, we have
E.10 Proof of Lemma C.3
Proof.
The proof of this lemma is inspired by [delanoe1991classical] and [urbas1997second]. Let be a point. There exists a neighborhood of such that is Then, by Assumption 3.2 and a version of Theorem 5.6 of [delfour1994shape], there exists a continuously differentiable function such that and for Notice that is an outward normal vector of at Let Also, there is such that To see this, recall that as is compact. Thus, there is such that As we have by Lemma E.3.
By the chain rule, we have
For a basis of the tangent space of at we have a decomposition
where the second equality follows because for any tangential vector on at Thus, it holds that
which implies that is normal to at Also, it is outward as ∎
E.11 Proof of Lemma C.4
Proof.
We fix such that satisfies condition (4). Notice that the set of such ’s has -measure one, as condition (4) holds almost surely. We omit arguments of functions if it does not make a confusion. Let Then, it holds that for
We consider the second term of the RHS. Since is a symmetric and positive definite matrix, we have the eigen decomposition where is an orthogonal matrix and is the diagonal matrix consisting of the eigenvalues. Let For any we have
where the inequality follows by Young’s inequality. In the same way, we also have, for any
Thus, it holds that
in which by setting
we obtain that the RHS is strictly positive unless by condition (4). ∎