### v4

# Local Identification in Instrumental Variable Multivariate Quantile Regression Models

Abstract
In the instrumental variable quantile regression (IVQR) model of [12], a one-dimensional unobserved rank variable monotonically determines a single potential outcome. In practice, when researchers are interested in multiple outcomes, it is common to estimate separate IVQR models for each of them. This approach implicitly assumes that the rank variable in each regression affects only its associated outcome, without influencing others. In reality, however, outcomes are often jointly determined by multiple latent factors, inducing structural correlations across equations.
To address this limitation, we propose a nonlinear instrumental variable model that accommodates multivariate unobserved heterogeneity, where each component of the latent vector acts as a rank variable corresponding to an observed outcome. When both the treatment and the instrument are discrete, we show that the structural function in our model is locally identified under a sufficiently strong positive correlation between the treatment and the instrument.
Keywords— IVQR, local identification, multivariate quantiles, optimal transport
1 Introduction
The instrumental variable quantile regression (IVQR) model introduced by [12] has become a widely used tool for estimating quantiles of potential outcomes in the presence of endogeneity. See, for example, [11] and [2]. While the IVQR framework assumes a one-dimensional outcome variable, researchers are often interested in settings with multiple outcomes. For instance, [11] examine the effect of 401(k) participation on several wealth measures, including total wealth and financial assets. In such cases, it is common to estimate separate IVQR models for each outcome dimension. However, this practice raises several conceptual and empirical concerns.
First, running separate models ignores the correlation structure among outcome variables and is therefore silent about their joint distribution. In the 401(k) example, this approach cannot capture the share of total wealth accounted for by financial assets.
More importantly, estimating separate quantile models makes rank similarity, a crucial assumption of [12], unrealistic. To apply the IVQR model in the 401(k) example, one must assume that individuals’ holdings of financial assets depend solely on their preferences for those assets. In reality, individuals allocate wealth across multiple assets simultaneously, and their preferences over financial assets are inherently linked to preferences over other asset categories.
To address these limitations, we propose a new nonlinear model that accommodates multidimensional outcomes that may be correlated with each other. Specifically, we consider a multivariate extension of the potential outcome framework in [12]. For a treatment , let denote the corresponding -dimensional potential outcome. Conditional on covariates, we assume it can be represented as , where is a structural function and is a -dimensional rank variable. The rank variable is a random vector that captures unobserved heterogeneity among observationally identical individuals, and can be interpreted as reflecting latent individual characteristics.
Existing studies of endogenous quantile models impose various monotonicity restrictions on the structural function for identification. Extending the assumption in [12] that the structural function is the quantile function of the potential outcome, we require its derivative to be symmetric and positive definite. This restriction implies that an increase in the th component of increases the th component of , while also allowing cross-dimensional effects: changes in one component of can affect other components of positively or negatively. Thus, our framework can capture substitutability and complementarity between outcome dimensions—features that the standard IVQR approach cannot accommodate. Section 2.2 presents two examples of structural functions satisfying this symmetry and positive definiteness condition.
Vector-valued functions with symmetric and positive definite derivatives play a central role in optimal transport theory, where they characterize maps minimizing quadratic transportation costs. Recent developments in statistics have shown that such maps possess many desirable properties that justify viewing them as multivariate analogues of quantile functions (see, e.g., [21], [9], [28], [26]). This perspective motivates us to refer to our framework as an instrumental variable multivariate quantile regression model.
For discrete treatments and instruments, our identification result generalizes the key insight of [12]. They show that to identify quantiles of one-dimensional potential outcomes under binary treatment, an instrument with binary or richer support satisfying the full-rank condition is required. For higher-dimensional potential outcomes, one might expect that instruments with more than two support points are necessary. Surprisingly, our main result, Theorem 3.1, shows that an instrument with only binary support suffices for identification. However, this comes at the cost of a stronger relevance condition: the instrument must be sufficiently positively correlated with the treatment, and the strength required increases with the dimension of the outcome vector.
Related literature. Our model builds on the IVQR framework proposed and developed by [12], [13], [15], and [14]. Unlike these studies, we allow for multidimensional outcomes and multidimensional unobserved heterogeneity.
Several related papers consider endogeneity in models with a single outcome variable. [1] analyze quantile treatment effects using the local average treatment effect framework of [29]. Their approach identifies treatment effects for compliers but is limited to binary treatments. Extending [16], [30] adopt a control function approach under the assumption that the selection equation is monotone and the instrumental variable is independent of the unobserved disturbance. Nonparametric identification in triangular models with discrete instruments is studied by [37], and further extended to multivariate settings by [27]. [31] investigates nonparametric identification in simultaneous equation models, where endogenous variables are continuous and their dimension must coincide with that of unobserved heterogeneity. In contrast, our framework accommodates discrete treatments and allows for arbitrary randomness in the treatment assignment process. Within the same simultaneous equation framework, [3] propose a nonseparable model that exploits proxy variables for unobserved heterogeneity.
Our key restriction on the structural function requires it to be a multivariate quantile function, as developed by [21], [9], [28], and [26]. Multivariate quantile functions are grounded in optimal transport theory ([39], [40]), which provides a natural framework for extending quantiles to random vectors. [25] offers a comprehensive overview of economic applications of optimal transport. In this paper, we draw on results from this literature to establish the identification of our model.
Organization: This paper is organized as follows. Section 2 introduces the IV multivariate quantile regression model. Section 3 formally states the local identification result for the model and discusses the relationship to existing results. Section 4 concludes. All proofs and mathematical preliminaries are provided in the Appendix.
2 IV multivariate quantile regression model
2.1 Model
Let be a -dimensional potential outcome vector, an endogenous variable (treatment), an exogenous covariate vector, and an IV. For now, the treatment and IV can be either discrete or continuous, although we will focus on the discrete cases for identification in the next section. Let be a compact convex set with a piecewisely boundary Let be a reference probability measure of which support is Assume is absolutely continuous with respect to the -dimensional Lebesgue measure.
Our model extends [12] so that it accounts for multidimensional outcome vectors. The following is the primitives.
Assumption 2.1.
The random variables satisfy the following conditions with probability one:
For each it holds that and there exists a function such that is continuously differentiable in the first variable, and its derivative is symmetric and positive definite on
For each is independent of conditional on
for some unknown function and random element
Conditional on are identically distributed.
The observed random variables consist of and
We refer to this framework as an IV multivariate quantile regression model. The term “multivariate quantile” follows the literature that extends the classical notion of quantile functions to multivariate random variables. Further details are provided in Appendix A. See also, for example, [6], [9], [28], and [26].
The difference from [12] appears in (A1). It follows from Theorem A.1 that for any reference measure there exists with a symmetric and positive semi-definite derivative such that but it is not necessarily regular as specified in (A1). Assumption (A1) assumes the smoothness of and the strict positive definiteness of its derivative. This assumption holds if both and the distribution of conditional on are sufficiently regular, according to the regularity theory of optimal transport (see, for example, [5]). Hence, for any other sufficiently regular reference measure condition (A1) is satisfied with some and This implies that any causal parameter defined as a functional of the distributions of the potential outcomes, such as average/quantile treatment effects, is independent of the choice of the reference measure. Also, the continuity of imposed in (A1) implies that the range of is compact and so is the support of conditional on
Condition (A2) implies that the IV is independent of potential outcomes conditional on covariates. Condition (A3) is a weak restriction that allows a broad class of assignment rules of treatments. In particular, in (A3) can be correlated with potential outcomes.
Condition (A4) is called the rank similarity ([12]). The simplest form of rank similarity is rank invariance that requires for all In the wealth accumulation example discussed in Section 1 and 2.2, the rank invariance assumes that an individual’s preference does not change regardless of their participation to 401(k). On the other hand, the rank similarity assumes that the preference may change depending on the participation status, but the individual cannot predict the change before deciding whether to participate 401(k) or not. For more details of these assumptions, see [12] and [14].
One limitation of the IV multivariate quantile regression model is that Condition (A4) depends on the choice of the reference measure . In particular, the fact that Condition (A4) holds for a given does not guarantee that it will hold for another measure . This dependence implies that the distribution of unobserved heterogeneity must be specified a priori. Although strong, such an assumption is common in related literature. For example, [10] impose a similar condition for identifying hedonic equilibrium models, and also, discrete-choice models typically rely on a fixed distribution for unobserved heterogeneity.
The unobserved random vector captures heterogeneity in outcomes among observationally identical individuals. It is interpreted as a rank variable in the sense that, conditional on covariates the th component of is monotonically determined by the corresponding component of In particular, under the positive definiteness of imposed in Assumption (A1), we have , where and . Moreover, unlike in the one-dimensional setting, the th component of may also influence the th component of for . Specifically, the cross-partial derivative may be either positive or negative, as long as remains positive definite. This feature enables the model to capture flexible correlations and potential substitutability or complementarity across different dimensions of the outcome vector.
Technically, any absolutely continuous probability measure on a compact convex set can be chosen as the reference measure . In the absence of a strong reason to prefer a particular specification, it is natural to take and , as in [6]. Alternatively, one may consider the -dimensional unit ball as and the spherical uniform distribution as , following [9] and [28].
By condition (A1), the observed outcome is written as the structural form where The following representation is the main testable implication of the model.
Theorem 2.1.
Suppose Assumption 2.1 holds. Then, it holds with probability one that for each measurable
where In particular, holds.
As the LHS of (1) is determined by the joint distribution of the observable variables, the equation gives a conditional moment restriction. In particular, any candidate structural function must satisfy (1) with replaced by . For identification, we explore conditions under which is the only function that satisfies the equation in the subsequent section.
For notational simplicity, let , suppressing the dependence on covariates . All subsequent analysis should be understood as conditional on . This simplification does not affect the identification results, since the conditional distribution of the observables given is identifiable.
When outcomes are univariate and the reference distribution is the uniform distribution on the unit interval, as in [12], is the -quantile of and satisfies
for all which corresponds to setting and in (1). If the treatment variable is supported on a finite set , this is a simultaneous equation with unknown variables Hence, if the instrument takes on more than or equal to values and the equation system is non-degenerate, it is expected that the solution is unique, at least locally. Indeed, [12] show that this observation is correct. Checking the condition for each quantile level establishes the nonparametric identification of the structural function
This approach works in one dimension because, for each , the values are determined by the conditional moment restriction (1) when we take the set . In higher dimensions, however, this argument no longer applies. To see why, consider the case with and fix . For any set , equation (1) does not isolate alone, because the image can be fully nonlinear. In particular, the image depends not only on the value of at but also on its behavior along the boundary . Since the set is an infinite-dimensional object, evaluating (1) for a fixed does not yield identification when the instrument has finite support. To address this issue, we instead interpret (1) as a measure-valued equation, allowing us to exploit the relationships between and for different sets and . Such relationships are unnecessary in the univariate case but are crucial for identifying multivariate structural functions.
2.2 Examples
Example 1. We first consider an extension of the saving model proposed by [11]. The paper investigates the effects of 401(k) participation on wealth accumulation. As measures of wealth, they consider total wealth, net financial assets, and net non-401(k) financial assets. Since [11] apply the IVQR model to these variables separately, their analysis does not capture the correlation among them. For example, while total wealth is defined as net financial assets plus other assets (e.g., housing equity and the value of business, property, and motor vehicles), their model has no implication for how people distribute total wealth into these sub categories. We shall see that our multivariate model can take both variables into consideration simultaneously and allows us to discuss their joint distribution.
Let represent the preference for net financial assets and other assets under participation status Let be net financial assets and other assets, respectively. An individual with type whose portfolio is receives a quasi-linear utility where is a deterministic concave function.111Notice that covariates are suppressed for simplicity. Assuming an interior solution, the individual chooses the optimal portfolio which follows by the first order condition. It can be seen that the function has a symmetric and positive definite derivative by the concavity of
The individual determines whether to participate 401(k) based on the eligibility of 401(k) and unobservable heterogeneity This is represented by The eligibility variable is considered to work as a valid IV. The joint distribution of fits into the multivariate IVQR model under and the rank invariance/similarity.
With identified, we can obtain the joint distribution of net financial assets and other assets. This allows us to discuss the effect of 401(k) participation on the portfolio selection, which is impossible in the standard IVQR framework.
Example 2. We now show how a simple discrete choice problem fits naturally into the IV multivariate quantile framework. This example is inspired by [35] and [24]. Consider a retailer offering differentiated goods in two sets of stores, indexed by The products’ qualities are captured by a latent quality vector where is assumed known (e.g., from in-store audits of freshness and display quality). Each consumer in district has an unobservable preference vector The preference distributions are unknown and may differ because district is exposed to a marketing campaign. By choosing product a consumer receives utility
The analyst never observes or . Instead, one observes the aggregate market share vector if the store is in district where
Economically, represents how a marginal improvement in the quality redistributes sales across all goods under the taste regime .
The Williams-Daly-Zachary theorem implies that the surplus function defined as
is related to the market share vector via the equation (see, e.g., [33]). It is not hard to see that is a convex function. Also, it is differentiable under mild regularity on Hence, has a symmetric and positive definite derivative.
Let denote the district, or equivalently the treatment status, ( or ) and suppose there exists a valid instrument (e.g. randomized ad‐exposure eligibility). Then we observe i.i.d. draws of with and This setup satisfies the rank invariance and exogeneity conditions of the IV multivariate quantile model.
Applying a separate IVQR mode of [12] to each product discards the rich cross‐product relationships. For instance, one-dimensional approaches cannot impose the natural constraint nor capture correlations in arising from similar products facing related quality shocks. The multivariate model preserves these economically meaningful cross‐product patterns while identifying the two demand-quality mappings and
2.3 Comparison with [12]
Even when multiple outcome variables are present, one might consider applying the one-dimensional quantile model of [12] to each component separately. However, this approach typically leads to the violation of the rank similarity assumption of [12]. To see this, we revisit the wealth accumulation example of Section 2.2. For consider a binary treatment environment and assume that satisfies Assumption 2.1 with the rank invariance and where is a random variable independent of all the other variables. Recall that is the vector consisting of net financial assets and other assets, is the participation status of 401(k), is its eligibility, and is the preference for the corresponding assets. The potential outcome for treatment is componentwisely represented as
Since and are assumed independent, it is expected that the treatment effect of is identified if is sufficiently informative.
Consider an empirical researcher who is interested only in net financial assets In this case, it is common to apply the standard quantile model to the data as [11] do. To see that this practice leads to the violation of the rank similarity assumption, notice that the structural equation under the one-dimensional model is where and is the quantile function of Then, is measurable in since The rank similarity requires but this is satisfied only when which does not hold in general.222For a counterexample, consider and In other words, the instrument is invalid in the sense that it is correlated with the unobserved heterogeneity via the treatment variable
The failure of rank similarity arises because depends not only on but also on whereas [12] do not allow such dependence—an assumption that is often unrealistic. Specifically, applying the framework of [12] to the wealth accumulation example would require assuming that individuals determine their holdings of financial assets solely based on their preference for financial assets. In reality, individuals also consider their preference for other assets, since it affects how much they can allocate to financial assets. Our multivariate quantile model accommodates such interdependence by allowing the potential outcomes to depend jointly on multiple unobservables thereby resolving this limitation.
In summary, the IV multivariate quantile model enables to consider multiple outcome variables at the same time, since it captures the correlation between different outcomes. Moreover, even if only some of the outcome variables are of interest, our model can alleviate the endogeneity by considering other outcomes together.
3 Identification
In this section, we consider the identification problem of the structural functions. We assume that the treatment and IV are supported on the same finite set, i.e., We will further restrict our attention to binary treatments/IVs in the later part of this section. More general cases are discussed in Appendix D.
The representation (1) of Theorem 2.1 and the change-of-variables formula imply that the true structural function solves the measure-valued equation
where
For a scalar function a vector-valued function and a matrix-valued function we define three types of supremum norms as follows:
For a compact set in a Euclidean space, let be the set of -valued functions on that are -times continuously differentiable on and the derivatives can be continuously extended to
Our identification result below shows that is the locally unique solution to (2) in a certain function class. To state the result formally, consider the normed space equipped with Also, for fixed constants define a subset of as
where is the largest eigenvalue of and define similarly. For a constant the space in which the parameter is identified is
Since we are interested in identification, we assume the correct specification.
Assumption 3.1.
Assumption 3.1 is weak under Assumption 2.1, as it just requires that be smooth and that it have moderate derivatives.
The conditions in the definition of restrict the class of admissible deviations from the true structural function. Such restrictions are common in the literature on nonparametric IV identification. To see their role, note that identifying requires inverting the linearization of the system of equations (2). Although this linearized system is invertible under the standard full-rank condition, its inverse is generally discontinuous because of the infinite-dimensional nature of the problem, rendering the identification ill-posed. This difficulty is resolved by restricting the identification domain to thereby ruling out pathological deviations from the truth. Similar regularity conditions are imposed in related work, including [8] and [7]. In particular, Section 2.3 of [8] presents an example where identification fails once an analogous restriction is removed.
We further impose two regularity conditions. Let be the support of the distribution of
Assumption 3.2.
For each the support is a convex compact set with a piecewisely boundary
Assumption 3.3.
For each
Assumption 3.2 requires that the support of a potential outcome be regular. The convexity of the support is necessary for many purposes in the optimal transport theory (see, for example, [39] and [23]). The smoothness of the boundary is rarely a problem in practice, as it allows for kinks at some points. Combined with Assumption 3.2, Assumption 3.3 implies that is bounded.
We consider the identification of structural functions in the following sense.
Definition 3.1.
Suppose We say the structural function is identified in if the following holds: if a set of random variable satisfies Assumption 2.1 with some then the joint distribution of is different from that of Also, we say the structural function is locally identified in if there exists such that is identified in
In what follows in this section, we focus on the case of i.e., to develop the idea clearly. Identification results for nonbinary treatments are given in Appendix D.
The following is our main theorem.
Theorem 3.1.
Condition (3) is a novel assumption that, to the best of our knowledge, has not appeared in the existing literature. Roughly speaking, it requires that () be sufficiently positively correlated with (). To see this, let denote the density of conditional on and and suppose it satisfies for some constants Then, the condition
is sufficient for condition (3) to hold. This inequality is satisfied when and are sufficiently large. Since condition (3) requires a positive correlation between and the labeling of the instrumental variable can be reversed if necessary to meet the condition.
Condition (3) is stronger than [12]’s monotone likelihood ratio condition:
This can be shown as follows. Since equation (3) implies
When condition (3) can be relaxed as follows.
Proposition 3.1.
Proposition 3.1 is consistent with [12], who show that identification holds when matrix (4) is full rank, since the determinant of any full-rank matrix can be made positive by switching its columns if necessary.
Identification for holds under a milder assumption because the proof of Proposition 3.1 relies on a property that is valid only in one dimension. Specifically, it uses the fact that the cofactor matrix333For an invertible matrix , its cofactor matrix is defined as . of any matrix equals the identity if and only if . See Lemma C.4 for details. This observation implies that replacing the positive definiteness of matrix (4) with condition (3) can be interpreted as the additional cost of allowing for multidimensional potential outcomes.
As the dimension increases, condition (3) becomes more demanding, which is natural given the need to identify higher-dimensional structural functions. However, the minimum support size of the instrumental variable required for identification remains independent of the dimension of the potential outcomes, as long as the IV is sufficiently correlated with the treatment variable.
3.1 Outline of the proof of Theorem 3.1
Roughly speaking, the proof of Theorem 3.1 proceeds in two steps. In the first step, we show that the support of the potential outcome is identified under condition (3). If a candidate structural function implies a different support for the potential outcome, then it cannot be consistent with the joint distribution of the observable variables Therefore, the identification analysis can be restricted to the class of structural functions that preserve the support.
In the second step, we show that is the locally unique solution to the system (2) of equations within the class of structural functions. The argument follows the logic of the implicit function theorem, which is commonly used in the identification analysis of nonlinear models with finite-dimensional parameters. Specifically, we linearize the nonlinear system (2) around the true structural function The key insight is that the positive correlation condition (3) ensures that the “slope” of the linearized system is bounded away from zero, thereby guaranteeing the local uniqueness of the solution.
4 Concluding Remarks
In this paper, we proposed a new nonlinear IV model that extends [12] to accommodate the correlation among multiple outcome variables. A key identifying restriction is that the structural functions are assumed to have a symmetric and positive definite derivative with respect to the rank vector. We showed that if the instrumental variable is sufficiently positively correlated with the treatment variable, the structural functions are locally identified. The cost of multidimensionality is that the positive correlation condition becomes more demanding as the dimension of the potential outcomes increases. Nevertheless, the minimum support size of the instrumental variable required for identification remains the same as in the univariate case. These results clarify how identification in nonlinear IV models can be extended to multidimensional settings while maintaining a similar structure to the scalar case.
Appendix A Preliminaries
Our multivariate outcome model is tightly related to the optimal transport theory. In this section, we briefly review mathematical concepts that are used in the paper. Fix a probability measure on a compact convex subset of the -dimensional Euclidean space, and let be a -dimensional random vector. First, consider the case of and According to the inverse probability integral transform, there exists a random variable such that almost surely where is the quantile function of (see, for example, Proposition 3.2 of [36]). Moreover, is the (almost surely) unique nondecreasing function that satisfies this equation for some uniform random variable. This observation for the classical one-dimensional case holds more generally by replacing the monotonicity of with the restriction that it is written as the gradient of a convex function.
Theorem A.1.
Let be an absolutely continuous probability measure on and be an absolutely continuous -dimensional random vector. There exists a convex function and a random vector such that almost surely and Furthermore, is -almost surely unique.
This is a version of the fundamental theorem in the optimal transport theory known as the Brenier-McCann theorem ([4] and [32]). The gradient is called the optimal transport map from to the distribution of under the quadratic cost because it solves the following Monge problem:
where is the Euclidean norm and is the distribution of
Given the reference probability measure the distribution of is characterized by the gradient via The function is called the multivariate quantile function of The multivariate quantile function has a symmetric and positive semi-definite derivative and therefore, it is cyclically monotone, that is, for any with it holds that
as is shown in [34]. Conversely, if a function has a symmetric and positive definite derivative, then it is written as the gradient of a convex function.
Lemma A.1.
Suppose that a differentiable function has a symmetric and positive definite derivative on Then, there exists a strictly convex differentiable function such that on Furthermore, is unique up to an additive constant.
Furthermore, a multivariate quantile function induces a bijection between the interior of the domain and that of the range under some regularity conditions. For similar results, see Theorem 1 of [17] and Proposition 3.1 of [26].
Theorem A.2.
Suppose that a continuous function has a symmetric and positive definite derivative on Also, assume that its image is convex. Then, is a continuous bijection between and
Appendix B Local identification for system of equations
B.1 Setup and theorem
The proof of Theorem 3.1 relies on a more general identification result (Theorem B.1), which is of independent interest. We devote this section to the identification result that can be applied to parameters that are characterized by the solution of a system of equations.
Let be normed spaces, be any subset, and be a finite set. For fix and We consider the following system of equations
Suppose is a solution to (6). We are interested in whether is the locally unique solution. To formulate deviations from we define the tangent set at as
The set specifies possible deviations in from Notice that if then We also define the set of normalized tangent vectors as
Remember that when and are finite-dimensional, the locally unique solvability of a system of equations is given by the implicit function theorem, which assumes the differentiability and the full rankness of the Jacobian matrix. As our identification result is an analogue statement for infinite-dimensional parameters, we impose assumptions corresponding to these two.
The first assumption is the differentiability of at
Assumption B.1.
For there exists such that
where if the set over which the supremum is taken is empty, the value reads zero.
It is clear that if Assumption B.1 is reduced to the standard Fréchet differentiability. For strict subsets of it is weaker than the Fréchet differentiability. In particular, even if is not a vector space, Assumption B.1 is well-defined. Unlike the Fréchet differentiability, the linearity of is not required.
The next assumption is an infinite-dimensional version of the full rankness of the Jacobian matrix.
Assumption B.2.
It holds that
Now, we are ready to state the main theorem in this section.
Theorem B.1.
This result is closely related to Section 2 of [8]. Indeed, their Theorems 1 can be shown from Theorem B.1. Our theorem has a different focus than their result in that it covers the local identification problem for systems of equations and that the domain of functionals is not assumed to be a Banach space.
B.2 Proof of Theorem B.1
We use the following lemma that gives an equivalent representation of Assumption B.2.
Lemma B.1.
Let Assumption B.2 holds if and only if there exists such that for all
Proof of Theorem B.1.
Take in Lemma B.1. By Assumption B.1, there is such that for all and
Fix and suppose that with satisfies for We will show By Assumption B.1,
where the equality holds because the first inequality holds because and the last one holds because By Assumption B.2 and Lemma B.1, Therefore,
which immediately implies ∎
Appendix C Proof of Theorem 3.1
Proof of Theorem 3.1.
We first show that the range of any candidate of the structural function must be the same as that of the truth which is
Lemma C.1.
Let Under condition (3), it holds
Since the set in the RHS of the statement is identified, if a candidate has a different range than the joint distribution of in the notation in Definition 3.1, cannot be the same as that of Thus, it is enough to show the local identification in
where
Next, we show that is the locally unique solution to the system (2) of equations in using Theorem B.1. Recall that is a normed space and that Let be the space of finite signed Borel measures on equipped with the total variation norm:
Then, is a normed space. For define as
Then, the identification restriction (2) is written as for all We apply Theorem B.1 to this system of equations where and The normalized tangent set that appears in Appendix B is
We need to check Assumptions B.1 and B.2. Remember that for a smooth vector field the divergence of at is defined as
The following lemma states that is differentiable in the sense of Assumption B.1.
Next, we show the Assumption B.2 by contradiction. Suppose there is a sequence such that
as By definition, there exists such that As and are optimal transport maps, they are written as and for some convex functions and By adding a constant, we may assume without loss of generality. Then satisfies and Since there exists such that it holds that for any
where Hence, we have
where the last inequality holds because
Let As we have which is independent of Also, it holds that and that
By considering a function in the definition (7) of the total variation norm, the integration by parts implies
where
where is the outward normal unit vector of at and is the -dimensional surface measure on
The folloiwng lemma shows that is nonnegative. Remember that since is continuous, implies
Lemma C.3.
For and define Then, at any point is an outward normal vector of
By Lemma C.3, we have
Since the support is convex by Assumption 3.2, it holds that for which implies Thus, we have and therefore,
Taking the sum over yields
Now, we restrict our attention to the case of The following lemma shows the positive definiteness of the integrand of the RHS of (9).
Lemma C.4.
By Lemma C.4, the integrand in the RHS of (9) is almost surely nonngetive. Since the LHS of (9) converges to zero by (8) and we have
for almost all by taking a subsequence if necessary. Recall that and since Then, the sequence is uniformly bounded and uniformly equicontinuous, and therefore, there exists a uniformly converging subsequence by the Arzelà–Ascoli theorem. Let be the limit function. By equation (10), almost everywhere by Lemma C.4, and it holds everywhere, as is continuous. Hence, the uniform convergence implies up to a subsequence. However, this contradicts to the fact that which particularly implies Thus, Assumption B.2 is satisfied.
Appendix D Generalization of Theorem 3.1
In this section, we state a generalization of Theorem 3.1 that allows for nonbinary treatments and IVs. The proof is a straightforward extension of that of Theorem 3.1 once we replace condition (3) with the following.
Assumption D.1.
There is a constant matrix such that for
holds for -almost all
Proof.
The proof is similar to that of Theorem 3.1. Lemma C.1 holds similarly under Assumption D.1, and Assumption B.1 holds as in Lemma C.2. We will verify Assumption B.2. For let
Clearly, Also, let Then, we have and Since
where it holds Notice also that Hence, by considering a function we have
where the last inequality follows by Lemma C.3. Moreover, by and Assumption D.1, it holds
The convergence of the LHS and Assumption D.1 imply that the integrand in the RHS converges to zero almost everywhere on up to a subsequence. Assumption D.1 also implies almost everywhere up to a subsequence. The Arzelà-Ascoli argument concludes which is a contradiction. ∎
Appendix E Omitted proofs
E.1 Proof of Theorem 2.1
E.2 Proof of Proposition 3.1
E.3 Proof of Theorem A.1
Proof.
By the Brenier-McCann theorem, there exists a convex function on such that where is the distribution of Let Then, it holds that Let be the Legendre transform of that is, Clearly, is convex. As for -almost all we have almost surely. The uniqueness of is a direct consequence of the Brenier-McCann theorem. ∎
E.4 Proof of Lemma A.1
Proof.
As is symmetric on which is a convex domain, there exists such that by Poincaré’s lemma. Since is positive definite, is strictly convex. The uniqueness follows immediately from Lemma 2.1 of [18]. ∎
E.5 Proof of Theorem A.2
Recall from the statement of Theorem A.2 that is continuous, surjective, and has a symmetric and positive definite derivative on We first provide three auxiliary lemmas.
Lemma E.1.
The function is strictly cyclically monotone on i.e., it is cyclically monotone, and the inequality of (5) is strict unless
Proof.
Lemma E.2.
The function is injective on
Proof.
Let be a convex function that satisfies the condition in Lemma A.1 for The strict convexity of implies that is injective on because its subdifferentials are disjoint. ∎
Lemma E.3.
For it holds that
Proof.
The proof is by contradiction. Suppose there is such that Since is convex, there exists such that for by the separating hyperplane theorem. It follows that for sufficiently small By Lemma E.1, we have Take a sequence such that Then, a contradiction. ∎
E.6 Proof of Lemma B.1
Proof.
The “if” part holds because for any
To show the “only if” part, suppose
where by the hypothesis. Then, we have
E.7 Proof of Lemma C.1
Proof.
Let such that for some Let Since we have where is the -ball around Thus, we have where the first equality holds by Assumption (A2). This implies and “” holds by taking the closure.
Let and suppose is not in the RHS. Then, there is such that for all and However, this implies that condition (3) fails on the ball, which has a positive measure. Hence, lies in the RHS, and “” holds by taking the closure. ∎
E.8 Proof of Lemma C.2
We first show an auxiliary lemma on the uniform differentiablity.
Lemma E.4.
Let be a subset of an inner product space Suppose the function is (Fréchet) differentiable on the interior of with a uniformly continuous derivative Then, is uniformly differentiable, i.e.,
Proof.
Let Since is uniformly continuous, there is such that
Fix and choose such that and Since
we have
Remark E.1.
In general, the integral of should be understood in the sense of Bochner, but this complication is not important for the proof of Lemma C.2.
For matrices let
For an invertible matrix define its cofactor matrix as
We rewrite the formula in Lemma C.2.
Lemma E.5.
For and it holds
Proof.
For simplicity of notation, we omit arguments of functions.
By the definition of the cofactor matrix, we have
For the second term, it holds
where the last equality follows from Piola’s identity. See, for example, the lemma in page 440 of [22]. ∎
Proof of Lemma C.2.
By Lemma E.4, we have
for small because by the definition of and because is continuously differentiable on by Assumption 3.3. Thus, it holds that
for small
Similarly, by Lemma E.4 and Jacobi’s formula, we have
for small because the eigenvalues of lie between and and because is of on This gives a bound for as follows:
for small as by the definition of
Finally, we have
E.9 Proof of Lemma C.3
Proof.
The proof of this lemma is inspired by [19] and [38]. Let be a point. There exists a neighborhood of such that is Then, by Assumption 3.2 and a version of Theorem 5.6 of [20], there exists a continuously differentiable function such that and for Notice that is an outward normal vector of at Let Also, there is such that To see this, recall that as is compact. Thus, there is such that As we have by Lemma E.3.
By the chain rule, we have
For a basis of the tangent space of at we have a decomposition
where the second equality follows because for any tangential vector on at Thus, it holds that
which implies that is normal to at Also, it is outward as ∎
E.10 Proof of Lemma C.4
Proof.
We fix such that satisfies condition (3). Notice that the set of such ’s has -measure one, as condition (3) holds almost surely. We omit arguments of functions if it does not make a confusion. Let Then, it holds that for
We consider the second term of the RHS. Since is a symmetric and positive definite matrix, we have the eigen decomposition where is an orthogonal matrix and is the diagonal matrix consisting of the eigenvalues. Let For any we have
where the inequality follows by Young’s inequality. In the same way, we also have, for any
Thus, it holds that
in which by setting
we obtain that the RHS is strictly positive unless by condition (3). ∎